# The Fractional-Logarithmic Laplacian: Fundamental Properties and Eigenvalues

### v1

@@ Abstract.
In this paper, we introduce, for the first time, the fractional–logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$, defined as the derivative of the fractional Laplacian $(-\Delta)^{t}$ at $t=s$. It is a singular integral operator with Fourier symbol $|\xi|^{2s}(2\ln|\xi|)$ and we prove the pointwise integral representation
@@
We develop the associated functional framework on both $\mathbb{R}^{n}$ and bounded Lipschitz domains, introducing the natural energy spaces and proving embedding results; in particular, we obtain a compact embedding at the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$, a phenomenon that differs from the classical Sobolev and fractional Sobolev settings. We further study the Poisson problem, proving existence and $L^{\infty}$-regularity results. We then investigate the Dirichlet eigenvalue problem and establish qualitative spectral properties. Finally, we derive a Weyl-type asymptotic law for the eigenvalue counting function and for the $k$-th Dirichlet eigenvalue, showing that the high-frequency behaviour combines the fractional Weyl scaling with the logarithmic growth factor, thus interpolating between the fractional Laplacian and the logarithmic Laplacian.
@@ 1. Introduction
is chosen so that $(-\Delta)^{s}$ has Fourier symbol $|\xi|^{2s}$ for $\xi\in\mathbb{R}^{n}$ (cf. [11, Chapter 3]), where $\Gamma$ denotes the Gamma function. With this normalization one also has the small-order limit
@@
Moreover, they proved (see [7, Theorem 1.1]) that $\operatorname*{Log}(-\Delta)$ has the (weakly) singular Fourier symbol $2\ln|\xi|$ for $\xi\in\mathbb{R}^{n}$, and admits the integral representation
@@
$L^{1}_{0}(\mathbb{R}^{n}):=\Bigl\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{n})\;:\;\int_{\mathbb{R}^{n}}(1+|x|)^{-n}|u(x)|\,dx<\infty\Bigr\}.$
Here $B_{1}(0)$ denotes the open unit ball in $\mathbb{R}^{n}$ centred at the origin and
$c_{n}=\frac{\Gamma(n/2)}{\pi^{n/2}}=\frac{2}{\omega_{n}},\qquad\rho_{n}:=2\ln 2+\psi(n/2)-\gamma,$
@@
More recently, the first and third authors, together with Weth [6], established that $\operatorname*{Log}(-\Delta)$ admits an extension property which is quite different from the Caffarelli-Silvestre extension for the fractional Laplacian (cf. [5]). Specifically, they showed that for every $u\in L^{1}_{0}(\mathbb{R}^{n})$,
@@
For every $0<s<1$, we define the operator $(-\Delta)^{s+\operatorname*{Log}}:C_{c}^{2}(\mathbb{R}^{n})\to\mathbb{\mathbb{R}}$ by
@@
for every $u\in C_{c}^{2}(\mathbb{R}^{n})$ and call $(-\Delta)^{s+\operatorname*{Log}}$ the fractional-logarithmic Laplacian.
@@
for $u\in L^{1}(\mathbb{R}^{n})$. We also denote the Lizorkin space (see [24, Section 5.1.2])
$\mathcal{Z}(\mathbb{R}^{n}):=\Bigl\{\varphi\in\mathscr{S}(\mathbb{R}^{n}):\ \partial^{\alpha}\widehat{\varphi}(0)=0,\ \text{for all }\alpha\in\mathbb{N}_{0}^{\,n}\Bigr\},$
as well as the corresponding topological dual $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$, i.e., the space of all continuous linear functionals on $\mathcal{Z}(\mathbb{R}^{n})$, where $\mathscr{S}(\mathbb{R}^{n})$ and $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ denote, respectively, the Schwartz space on $\mathbb{R}^{n}$ and the space of tempered distributions on $\mathbb{R}^{n}$. With this choice, for every symbol $m(\xi)$ in
$m(\xi)\in\Bigl\{|\xi|^{2s},\ \ln|\xi|^{2},\ |\xi|^{2s}\ln|\xi|^{2}\Bigr\},$
the associated operator $m(D)$ defines a continuous linear map $\mathcal{Z}^{\prime}(\mathbb{R}^{n})\to\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ by duality,
$\langle m(D)u,\varphi\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}:=\Big\langle u,\ \mathcal{F}^{-1}\!\big(m(\xi)\,\widehat{\varphi}(\xi)\big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}},\qquad u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n}),\ \varphi\in\mathcal{Z}(\mathbb{R}^{n}).$
In particular, $(-\Delta)^{s}$, $\operatorname*{Log}(-\Delta)$ and $(-\Delta)^{s+\operatorname*{Log}}$ are well-defined continuous linear operators on $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$.
@@
    (Representation as a composition) For every $u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n})$
    $(-\Delta)^{s+\operatorname*{Log}}u=(-\Delta)^{s}\bigl(\operatorname*{Log}(-\Delta)u\bigr)=\operatorname*{Log}(-\Delta)\bigl((-\Delta)^{s}u\bigr)\quad\text{in}\quad\mathcal{Z}^{\prime}(\mathbb{R}^{n}).$
@@
Concerning Theorem 1.1, we record the following remark.
@@
By the spectral theorem, we can define the fractional Laplacian
$(-\Delta)^{s}=\int_{[0,\infty)}\lambda^{s}\,dE(\lambda):H^{2s}(\mathbb{R}^{n})\rightarrow L^{2}(\mathbb{R}^{n})$
with domain (spectral fractional space)
$H^{2s}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{[0,\infty)}\lambda^{2s}\,dE_{u,u}(\lambda)<\infty\Bigr\},$
see Section 2.2 for more details. Since
@@
$\bigl\langle(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}u,v\bigr\rangle_{L^{2}}=\int_{(0,\infty)}\lambda^{s}\ln\lambda\,dE_{u,v}(\lambda),\quad u\in\operatorname{Dom}((-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}),\ v\in L^{2}(\mathbb{R}^{n}),$
where
(1.8) $\operatorname{Dom}((-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}})=\Bigl\{u\in L^{2}(\mathbb{R}^{n}):\int_{(0,\infty)}\lambda^{2s}\ln^{2}\lambda\,dE_{u,u}(\lambda)<\infty\Bigr\}=:H^{2s+\operatorname*{Log}}(\mathbb{R}^{n}).$
@@
Therefore, by Proposition 1.2 and the definition of fractional-logarithmic Laplacian, we can obtain that the spectral operator $(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}$ coincides with $(-\Delta)^{s+\operatorname*{Log}}$ in $C_{c}^{2}(\mathbb{R}^{n})$. By the relation between the spectral measure of $\Delta$ and the Fourier transform (cf. [23, Chapter 7]):
@@
Equipped with this norm, $\mathcal{H}^{s}_{0}(\Omega)$ is a Hilbert space. It is well known that (see [18, Part I])
@@ 2.1. Equivalent Represnetations
(iii) It follows by [11, Chapter 3] that
@@ 3.1. Properties of Functional Spaces
Set z=xyz=x-y and arguing as in the standard fractional Sobolev case, see [11, Proposition 3.4] one gets
@@ 4. The Dirichlet eigenvalue problem
is weakly lower semicontinuous (in fact, it is even of the class $C^{1}$). Thus, the remaining proof of Theorem 1.5 is similar to the one of [7, Theorem 1.4]. For the convenience of the readers, we provide here the details.
@@
It follows by [14, Theorem 2] that
@@
From [17, Theorem 12.8], there holds
@@
Similar to the proof of [14, (8)], there exists $C>0$ such that

### v3

@@ Abstract.
In this paper, we introduce the fractional–logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$, defined as the derivative of the fractional Laplacian $(-\Delta)^{t}$ at $t=s$. It is a singular integral operator with Fourier symbol $|\xi|^{2s}(2\ln|\xi|)$ and we prove the pointwise integral representation
@@
We develop the associated functional analytical framework for studying the fractional–logarithmic operator $(-\Delta)^{s+\operatorname*{Log}}$ on $\mathbb{R}^{n}$ and on bounded Lipschitz domains. We introduce the natural corresponding energy spaces and establish Sobolev and Poincaré embeddings; in particular, we obtain a compact embedding at the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$, a phenomenon that differs from the classical Sobolev and fractional Sobolev settings.
We further study the Poisson problem, proving existence and $L^{\infty}$-regularity results. We then investigate the Dirichlet eigenvalue problem and establish qualitative spectral properties. Finally, we derive a Weyl-type asymptotic law for the eigenvalue counting function and for the $k$-th Dirichlet eigenvalue, showing that the high-frequency behaviour combines the fractional Weyl scaling with the logarithmic growth factor, thus interpolating between the fractional Laplacian and the logarithmic Laplacian.
@@ 1. Introduction
is chosen so that $(-\Delta)^{s}$ has Fourier symbol $|\xi|^{2s}$ for $\xi\in\mathbb{R}^{n}$ (cf. [5, Chapter 3]), where $\Gamma$ denotes the Gamma function. With this normalization one also has the small-order limit
@@
Moreover, they proved (see [6, Theorem 1.1]) that $\operatorname*{Log}(-\Delta)$ has the (weakly) singular Fourier symbol $2\ln|\xi|$ for $\xi\in\mathbb{R}^{n}$, and admits the integral representation
@@
$L^{1}_{0}(\mathbb{R}^{n}):=\Bigl\{u\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{n})\;\Big|\;\int_{\mathbb{R}^{n}}(1+|x|)^{-n}|u(x)|\,dx<\infty\Bigr\}.$
Here, and throughout this paper, we write $B_{1}(0)$ to denote the open unit ball in $\mathbb{R}^{n}$ centred at the origin, and set
$c_{n}=\frac{\Gamma(n/2)}{\pi^{n/2}}=\frac{2}{\omega_{n}}\quad\text{ and }\quad\rho_{n}:=2\ln 2+\psi(n/2)-\gamma,$
@@
More recently, the first and third authors established in joint work with Weth [7] that the logarithmic Laplacian $\operatorname*{Log}(-\Delta)$ admits an extension property, and this one is quite different from the Caffarelli-Silvestre extension for the fractional Laplacian (see [8]). Specifically, they showed that for every $u\in L^{1}_{0}(\mathbb{R}^{n})$,
@@
Let $0<s<1$. Then, we call the operator $(-\Delta)^{s+\operatorname*{Log}}:C_{c}^{2}(\mathbb{R}^{n})\to\mathbb{\mathbb{R}}$ defined by
@@
for every $u\in C_{c}^{2}(\mathbb{R}^{n})$ the fractional-logarithmic Laplacian $(-\Delta)^{s+\operatorname*{Log}}$.
@@
for $u\in L^{1}(\mathbb{R}^{n})$. We also denote the Lizorkin space (see [18, Section 5.1.2])
$\mathcal{Z}(\mathbb{R}^{n}):=\Bigl\{\varphi\in\mathscr{S}(\mathbb{R}^{n})\,\big|\,\partial^{\alpha}\widehat{\varphi}(0)=0\ \text{for all }\alpha\in\mathbb{N}_{0}^{\,n}\Bigr\},$
as well as the corresponding topological dual $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$, i.e., the space of all continuous linear functionals on $\mathcal{Z}(\mathbb{R}^{n})$, where $\mathscr{S}(\mathbb{R}^{n})$ and $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ denote, respectively, the Schwartz space on $\mathbb{R}^{n}$ and the space of tempered distributions on $\mathbb{R}^{n}$. Within this framework, for every symbol $a(\xi)$ of the class
$\Bigl\{|\xi|^{2s},\ 2\,\ln|\xi|,\ |\xi|^{2s}\,2\,\ln|\xi|\Bigr\},$
the associated singular pseudo-differential operator $A(D)$ defines a continuous linear map $A(D):\mathcal{Z}^{\prime}(\mathbb{R}^{n})\to\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ given by (duality)
$\langle A(D)u,\varphi\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}:=\Big\langle u,\ \mathcal{F}^{-1}\!\big(a(\xi)\,\widehat{\varphi}(\xi)\big)\Big\rangle_{\mathcal{Z}^{\prime},\mathcal{Z}}$
for every $u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n})$ and $\varphi\in\mathcal{Z}(\mathbb{R}^{n})$. In particular, the operators $(-\Delta)^{s}$, $\operatorname*{Log}(-\Delta)$ and $(-\Delta)^{s+\operatorname*{Log}}$ are well-defined continuous linear operators on $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$.
@@
    (Representation as a composition) For every $u\in\mathcal{Z}^{\prime}(\mathbb{R}^{n})$,
    $(-\Delta)^{s+\operatorname*{Log}}u=(-\Delta)^{s}\bigl(\operatorname*{Log}(-\Delta)u\bigr)=\operatorname*{Log}(-\Delta)\bigl((-\Delta)^{s}u\bigr)$
    in $\mathcal{Z}^{\prime}(\mathbb{R}^{n})$.
@@
Concerning Theorem 1.1, we have the following remarks.
@@
By the spectral theorem, the fractional Laplacian $(-\Delta)^{s}:H^{2s}(\mathbb{R}^{n})\rightarrow L^{2}(\mathbb{R}^{n})$ is defined by
$(-\Delta)^{s}=\int_{[0,\infty)}\lambda^{s}\,dE(\lambda)$
with domain
$H^{2s}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n})\,\Big|\,\int_{[0,\infty)}\lambda^{2s}\,dE_{u,u}(\lambda)<\infty\Bigr\},$
which we call the spectral fractional Sobolev space. We refer to Section 2.2 for more details. Since
@@
$\bigl\langle(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}u,v\bigr\rangle_{L^{2}}=\int_{(0,\infty)}\lambda^{s}\ln\lambda\,\mathrm{d}E_{u,v}(\lambda)$
for every $u\in H^{2s+\operatorname*{Log}}(\mathbb{R}^{n})$, and $v\in L^{2}(\mathbb{R}^{n})$ with domain
(1.8) $H^{2s+\operatorname*{Log}}(\mathbb{R}^{n}):=\Bigl\{u\in L^{2}(\mathbb{R}^{n})\,\Big|\,\int_{(0,\infty)}\lambda^{2s}\ln^{2}\lambda\,dE_{u,u}(\lambda)<\infty\Bigr\}.$
@@
Therefore, by Proposition 1.2 and the definition of fractional-logarithmic Laplacian, we can obtain that the spectral operator $(-\Delta)_{\mathrm{spec}}^{s+\operatorname*{Log}}$ coincides with $(-\Delta)^{s+\operatorname*{Log}}$ in $C_{c}^{2}(\mathbb{R}^{n})$. By the relation between the spectral measure of $\Delta$ and the Fourier transform (cf. [20, Chapter 7]):
@@
Equipped with this norm, $\mathcal{H}^{s}_{0}(\Omega)$ is a Hilbert space. It is well known that (see [3, Part I])
@@ 2.1. Equivalent Representations
(iii) It follows by [5, Chapter 3] that
@@ 3.1. Properties of Functional Spaces
Set z=xyz=x-y and arguing as in the standard fractional Sobolev case, see [5, Proposition 3.4] one gets
@@ 4. The Dirichlet eigenvalue problem
is weakly lower semicontinuous (in fact, it is even of the class $C^{1}$). Thus, the remaining proof of Theorem 1.5 is similar to the one of [6, Theorem 1.4]. For the convenience of the readers, we provide here the details.
@@
It follows by [22, Theorem 2] that
@@
From [23, Theorem 12.8], there holds
@@
Similar to the proof of [22, (8)], there exists $C>0$ such that
@@
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.